A proton moves at 4.45 × 107 m/s perpendicular to a magnetic field. The field causes the proton to travel in a circular path of radius of 0.09 m. What is the field strength? (mproton = 1.6726 × 10-27 kgs and q=1.602×10−19 C)


1
Expert's answer
2021-05-24T09:02:45-0400

The proton experiences the action of Lorentz force:

"F = qvB"

At the same time, according to the second Newton's law, for a particle moving on the circular orbit

"\displaystyle F = ma = \frac{mv^2}{R}"

"\displaystyle \frac{mv^2}{R} = qvB"

"\displaystyle B=\frac{mv}{qR}= \frac{1.6726 \cdot 10^{-27} \cdot 4.45 \cdot 10^7}{1.602 \cdot 10^{-19} \cdot 0.09} =5.1623 \; [T]"


Answer: 5.1623 T